can someone help me solve this equation with steps?
-5x-3y=-10
-15x-7y=-10
it says to solve each system on the instructions.

Answer by Alan3354(69443)

-5x-3y=-10
-15x-7y=-10
-------------
Multiply the 1st eqn by -3, then add them.